# Dynamical Theory of Elastic Synchronization of Cardiomyocytes

### 15 Apr 2026 v1

@@ Abstract
where $S(\bm{r})$ is a Gaussian shape function defined by
$S(\bm{r})=\exp\!\left(-\frac{|\bm{r}|^{2}}{2\sigma_{\rm cell}^{2}}\right),$ (6)
@@
Figure 3: Orientation-dependent dynamical state maps for (a) $d/l_{\rm cell}=2.5$ and (b) $d/l_{\rm cell}=6$. Blue and red regions represent configurations yielding in-phase ($|\Delta\phi^{*}|<0.1$) and anti-phase ($|\pi-|\Delta\phi^{*}||<0.1$) synchronization, respectively. In these regions, the brightness indicates the coupling amplitude $\Delta\Gamma$, where lighter shades correspond to weaker interaction. Black regions represent configurations not classified as in-phase or anti-phase; they appear near the state boundary and are associated with very small coupling amplitude ($\Delta\Gamma<4\times 10^{-6}$).
@@
To construct the dynamical state maps (Fig. 3), we classified the stable phase differences $\Delta\phi^{\ast}$ obtained from phase reduction theory. Configurations were categorized as in-phase if $|\Delta\phi^{\ast}|<0.1$ (blue) and anti-phase if $|\pi-|\Delta\phi^{\ast}||<0.1$ (red). The intermediate states that did not satisfy either criterion are shown in black. The coupling amplitude $\Delta\Gamma=\Gamma_{\max}-\Gamma_{\min}$ varies continuously and becomes small near the state boundary, where the interaction is significantly weakened and the synchronization time becomes longer. In the intermediate states, we find $\Delta\Gamma<4\times 10^{-6}$ and $|\Delta\phi^{\ast}|$ distributed near 0 and $\pi$. Thus, they represent marginal states at the state boundary rather than distinct out-of-phase states. Comparing the results for $d/l_{\rm cell}=2.5$ and $d/l_{\rm cell}=6$, we find that the state boundary exhibits a similar shape in the $(\theta_{1},\theta_{2})$ plane, reflecting the robust orientation-dependence of the phase selection.
We next examine the distance dependence of the coupling strength. As shown in Fig. 4, the coupling amplitude $\Delta\Gamma$ decays as $d^{-3}$. This reflects the dipolar nature of the interaction and implies a long characteristic synchronization time, $\tau_{\rm sync}\sim 1/\Delta\Gamma\sim 10^{3}$$10^{5}$. Combined with the state maps in Fig. 3, this result suggests that the relative angular alignment primarily determines the selected synchronized state, while the spatial separation mainly affects the coupling strength and the synchronization time.
@@
For the representative configurations considered previously, our dynamical analysis agrees with the earlier energetic prediction [3]. This agreement is nevertheless nontrivial. Indeed, in the small-deformation approximation, the substrate-mediated elastic energy can be written as $H=\frac{1}{2}\bm{X}^{\rm T}\bm{M}^{-1}\bm{X}$, so that the reaction force is conservative, $\hat{F}_{i}=-\partial H/\partial X_{i}$. However, the synchronized state is selected through the non-isochronous part of the phase dynamics, $\dot{\phi}_{i}-\Omega_{0}=-Z_{2}(\phi_{i})\hat{F}_{i}$, where $Z_{2}$ denotes the second component of the phase sensitivity function. This component is not simply proportional to $\partial X/\partial\phi$ in general. Therefore, the phase velocity is not necessarily proportional to the energy gradient, $\partial H/\partial\phi_{i}=\hat{F}_{i}(\partial X_{i}/\partial\phi_{i})$. An analytical approach in simplified limits may further clarify the geometric structure of the coupling, but this lies beyond the scope of the present Letter and will be discussed elsewhere.

### 18 Apr 2026 v2

@@ Abstract
where $F_{i}(t)$ is the force exerted by each pole on the substrate, and $S(\bm{r})$ is a normalized Gaussian shape function defined by
$S(\bm{r})=\frac{1}{2\pi\sigma_{\rm cell}^{2}}\exp\!\left(-\frac{|\bm{r}|^{2}}{2\sigma_{\rm cell}^{2}}\right),$ (6)
@@
Figure 3: Orientation-dependent dynamical state maps for (a) $d/l_{\rm cell}=2.5$ and (b) $d/l_{\rm cell}=6$. Blue and red regions represent configurations yielding in-phase ($|\Delta\phi^{*}|<0.1$) and anti-phase ($|\pi-|\Delta\phi^{*}||<0.1$) synchronization, respectively. In these regions, the brightness indicates the coupling amplitude $\Delta\Gamma$, where lighter shades correspond to weaker interaction. Black regions represent configurations not classified as in-phase or anti-phase; they appear near the state boundary and are associated with small coupling amplitude ($\Delta\Gamma<7\times 10^{-4}$ ).
To construct the dynamical state maps (Fig. 3), we classified the stable phase differences $\Delta\phi^{\ast}$ obtained from phase reduction theory. Configurations were categorized as in-phase if $|\Delta\phi^{\ast}|<0.1$ (blue) and anti-phase if $|\pi-|\Delta\phi^{\ast}||<0.1$ (red). The intermediate states that did not satisfy either criterion are shown in black. The coupling amplitude $\Delta\Gamma=\Gamma_{\max}-\Gamma_{\min}$ varies continuously and becomes small near the state boundary, where the interaction is significantly weakened and the synchronization time becomes longer. In the intermediate states, we find $\Delta\Gamma<$ $7\times 10^{-4}$ and $|\Delta\phi^{\ast}|$ is often found near 0 and $\pi$. Thus, they represent marginal states at the state boundary rather than distinct out-of-phase states. Comparing the results for $d/l_{\rm cell}=2.5$ and $d/l_{\rm cell}=6$, we find that the state boundary exhibits a similar shape in the $(\theta_{1},\theta_{2})$ plane, reflecting the robust orientation-dependence of the phase selection.
[…]
@@
For the representative configurations considered previously, our dynamical analysis agrees with the earlier energetic prediction regarding the selection of in-phase or anti-phase synchronization. [3] This agreement is nevertheless nontrivial. Indeed, in the small-deformation approximation, the substrate-mediated elastic energy can be written as $H=\frac{1}{2}\bm{X}^{\rm T}\bm{M}^{-1}\bm{X}$, so that the reaction force is conservative, $\hat{F}_{i}=-\partial H/\partial X_{i}$. However, the synchronized state is selected through the non-isochronous part of the phase dynamics, $\dot{\phi}_{i}-\Omega_{0}=-Z_{2}(\phi_{i})\hat{F}_{i}$, where $Z_{2}$ denotes the second component of the phase sensitivity function. This component is not simply proportional to $\partial X/\partial\phi$ in general. Therefore, the phase velocity is not necessarily proportional to the energy gradient, $\partial H/\partial\phi_{i}=\hat{F}_{i}(\partial X_{i}/\partial\phi_{i})$. An analytical approach in simplified limits may further clarify the geometric structure of the coupling, but this lies beyond the scope of the present Letter and will be discussed elsewhere.